OpenStudy (anonymous):

The measure of an interior angle of a regular polygon is 170°. How many sides make up the regular polygon?

OpenStudy (anonymous):

\[\left( n-2 \right)*180/n\] gives you the size of the interior angle of any regular polygon with side n. this means that: \[\left( n-2 \right)*180/n\] = 170 solve for n \[\left( n-2)*180 \right) = 170n\] 180n -360 = 170n 10n =360 n = 36

OpenStudy (anonymous):

Interior 170 -> exterior = 10 360/10 = 36